### v1

@@ Abstract
This paper uses the three EHR datasets, MIMIC-IV (Johnson et al., 2023), eICU (Pollard et al., 2018), and HiRID (Hyland et al., 2020), which are publicly available on the PhysioNet repository (Johnson et al., 2020; Pollard et al., 2019; Faltys et al., 2021). More details about datasets can be found at Section 4.1. Our implementation code can be accessed at this repository.111https://anonymous.4open.science/r/LabTOP-DE7B
@@ 1 Introduction
Meanwhile, the advancement of machine learning has greatly contributed to healthcare for several clinical prediction tasks such as readmission, mortality risk, and length of hospital stay (Song et al., 2018; Xiao et al., 2018; Hur et al., 2022, 2023; Kim et al., 2023; Wornow et al., 2023; Renc et al., 2024). Although these works have demonstrated the value of lab test data in supporting clinical decisions, there has been no attempt to directly estimate various lab test results within a single unified model. Instead, most prior studies approach lab test outcome prediction by classifying discrete levels of a small subset of selected lab items (Hur et al., 2023; Kim et al., 2023; Wornow et al., 2023). Even among studies that attempt to estimate continuous values, they are specifically designed for only a few selected lab tests rather than whole set of lab items (Zhi et al., 2024; Jiang et al., 2024; Duan et al., 2020; Liu et al., 2023; Fu et al., 2023; Langarica et al., 2024). This narrow focus limits their practical applicability in clinical settings, where precise numerical predictions for a broader range of lab test results can facilitate early intervention and support more informed clinical decision-making across diverse medical conditions.
@@ 2 Related Works
Although previous studies have significantly advanced the field of machine learning for healthcare, none of them have addressed the direct estimation of continuous values for various lab tests within a single unified model. Instead, most of previous studies have explored the prediction of lab test results by designing a specific model for only a few selected lab items or classifying discrete levels of selected lab items, such as normal or abnormal ranges. For example, HgbNet (Zhi et al., 2024) employed an attention-based LSTM model to estimate hemoglobin levels, while Jiang et al. (2024) presented a probabilistic model using NGBoost (Duan et al., 2020) to predict expected values of several lab items, such as wbc and hemoglobin. Similarly, several works (Liu et al., 2023; Fu et al., 2023; Langarica et al., 2024) attempted to estimate glucose values using machine learning or deep learning approach. However, these approaches are specifically designed for a small subset of lab tests, making it challenging to generalize across diverse clinical measurements.
@@ 4.4 Results
tab:main MIMIC-IV eICU HiRID NMAE Naive Naive() GenHPF XGBoost GPT-4o GPT-4o-mini LabTOP SMAPE (%) Naive Naive() GenHPF XGBoost GPT-4o GPT-4o-mini LabTOP
@@
Interestingly, LabTOP also outperforms GPT-4o and GPT-4o-mini, confirming that general-purpose LLMs struggle with lab test outcome prediction task. This underscores the importance of explicitly learning EHR data rather than relying on general LLM capabilities. Furthermore, GPT-4o-mini exhibits significantly lower performance, which we attribute to its tendency to generate unrealistic lab test values. Upon manual inspection, we found that GPT-4o-mini often outputs numerically implausible values that are far outside the expected range (e.g., predicting ph of arterial blood as 135 whereas the ground-truth is 7.2). This result demonstrates that our data processing and training strategy play a crucial role in achieving robust performance by effectively leading the model to produce clinically meaningful predictions.
@@ 5 Conclusion
Our experiments have some limitations as follows. First, LabTOP requires relatively longer sequence length, leading to increased computational costs during training and inference, especially when we involve more types of EHR events. Future work could explore more efficient tokenization strategies to balance precision and computational efficiency. Second, our experiments were focused on retrospective EHR data, and its real-world applicability in prospective clinical settings remains to be validated. Further research could assess LabTOP’s performance in real-time hospital environments and evaluate its impact on clinical decision-making.
@@ Appendix B Implementation Details
To create samples for GPT-4o and GPT-4o-mini555Specifically, at the time of the experiments, GPT-4o referred to gpt-4o-2024-11-20 and GPT-4o-mini corresponded to gpt-4o-mini-2024-07-18., we followed the same strategy with GenHPF. In other words, we grouped test samples that have the same prior medical history (i.e., the lab items that occurred at the same time), and the LLM was instructed to answer values of the target lab items based on their prior sequence of medical events as prompt. The prompt we used for lab test outcome prediction can be found in Figure 3.

### v3

@@ Abstract
This paper uses the three EHR datasets, MIMIC-IV (Johnson et al., 2023), eICU (Pollard et al., 2018), and HiRID (Hyland et al., 2020), which are publicly available on the PhysioNet repository (Johnson et al., 2020; Pollard et al., 2019; Faltys et al., 2021). More details about datasets can be found at Section 4.1. Our implementation code can be accessed at this repository.111https://github.com/sujeongim/LabTOP
@@ 1 Introduction
Meanwhile, the advancement of machine learning has greatly contributed to healthcare for several clinical prediction tasks such as readmission, mortality risk, and length of hospital stay (Song et al., 2018; Xiao et al., 2018; Hur et al., 2022, 2023; Kim et al., 2023; Wornow et al., 2023; Renc et al., 2024). While previous studies have demonstrated the value of lab test data in supporting clinical decisions, efforts to estimate various lab test results within a single unified model have been limited. Notably, most prior studies approach lab test outcome prediction by classifying discrete levels of a small subset of selected lab items (Hur et al., 2023; Kim et al., 2023; Wornow et al., 2023). Even among studies that attempt to estimate continuous values, they are specifically designed for only a few selected lab tests rather than whole set of lab items (Zhi et al., 2024; Jiang et al., 2024; Duan et al., 2020; Liu et al., 2023; Fu et al., 2023; Langarica et al., 2024). On the other hand, some works on irregularly sampled time-series models have explored the prediction of multiple lab measurements using neural ODE-based approaches (Rubanova et al., 2019; Schirmer et al., 2022). These methods estimate multiple lab values within a unified framework; however, they primarily focus on modeling the dynamics of lab test results alone, without incorporating the broader medical events present in EHRs, such as medications and input events. This narrow focus limits their practical applicability in clinical settings, where precise numerical predictions for a broader range of lab test results can facilitate early intervention and support more informed clinical decision-making across diverse medical conditions.
@@ 2 Related Works
Although previous studies have significantly advanced the field of machine learning for healthcare, only a few of them have addressed the direct estimation of continuous values for various lab tests within a single unified model. For example, there have been attempts to estimate multiple lab values within a unified framework using neural ODE-based approaches (Rubanova et al., 2019; Schirmer et al., 2022). However, a key limitation of these approaches is that they only model lab test values in isolation, without incorporating other essential patient events such as medications, interventions, and clinical inputs that may significantly influence lab test dynamics. Instead, most of previous studies have explored the prediction of lab test results by designing a specific model for only a few selected lab items or classifying discrete levels of selected lab items, such as normal or abnormal ranges. For example, HgbNet (Zhi et al., 2024) employed an attention-based LSTM model to estimate hemoglobin levels, while Jiang et al. (2024) presented a probabilistic model using NGBoost (Duan et al., 2020) to predict expected values of several lab items, such as wbc and hemoglobin. Similarly, several works (Liu et al., 2023; Fu et al., 2023; Langarica et al., 2024) attempted to estimate glucose values using machine learning or deep learning approach. However, these approaches are specifically designed for a small subset of lab tests, making it challenging to generalize across diverse clinical measurements.
@@ 4.3 Baselines
LLaMA-3.1-Instruct (8B) (Grattafiori et al., 2024): We also evaluate the open-sourced instruction-tuned LLM designed for general-purpose tasks. For the experiments on this large model, we sample 10% of the test set as the same with GPT-4o and GPT-4o-mini.
@@ 4.4 Results
tab:main MIMIC-IV eICU HiRID NMAE Naive Naive() GenHPF XGBoost GPT-4o GPT-4o-mini LLaMA-3.1-Inst. LabTOP SMAPE (%) Naive Naive() GenHPF XGBoost GPT-4o GPT-4o-mini LLaMA-3.1-Inst. LabTOP
@@
Interestingly, LabTOP also outperforms general-purpose LLMs (GPT-4o, GPT-4o-mini, and LLaMA-3.1-Instruct), confirming that these models struggle with lab test outcome prediction task. This underscores the importance of explicitly learning EHR data rather than relying on general LLM capabilities. Furthermore, GPT-4o-mini and LLaMA-3.1-Instruct exhibit significantly lower performance, which we attribute to its tendency to generate unrealistic lab test values. Upon manual inspection, we found that GPT-4o-mini and LLaMA-3.1-Instruct often output numerically implausible values that are far outside the expected range (e.g., predicting ph of arterial blood as 135 whereas the ground-truth is 7.2). This result demonstrates that our data processing and training strategy play a crucial role in achieving robust performance by effectively leading the model to produce clinically meaningful predictions.
@@ 5 Conclusion and Discussion
Clinical utility of LabTOP
Our model allows for the quantification of uncertainty in its predictions, which can be measured using the average entropy of token probabilities when generating lab test values. For example, if a model outputs “1 2 3 . 4 5”, we can compute entropies for the generated tokens (“1”, “2”, “3”, “.”, “4”, and “5”) and average them to get the uncertainty of the predicted output. By leveraging this uncertainty estimation, high-confidence predictions can be directly used for clinical decision support, while low-confidence predictions can be flagged for further verification through actual lab testing. This approach helps prioritize critical tests while reducing unnecessary routine tests. As a specific scenario, in ICUs and other hospital settings where continuous monitoring is critical, patients frequently undergo multiple daily lab tests to assess their condition. However, frequent blood draws can lead to increased risk of anemia and additional workload for healthcare providers. According to James et al. (2018), ICU patients often undergo 1-4 blood tests per day, and repeated phlebotomy is a major contributor to anemia in critically ill patients. In this scenario, by incorporating LabTOP into clinical workflows, healthcare providers could pre-screen lab test predictions and selectively order confirmatory tests only when predictions exhibit high uncertainty. This approach would reduce unnecessary testing, helping to minimize patient burden and optimize resource allocation while maintaining diagnostic accuracy.
@@
Our experiments have some limitations as follows. First, LabTOP requires relatively longer sequence length, leading to increased computational costs during training and inference, especially when we involve more types of EHR events. To mitigate this, we applied optimization techniques such as mixed-precision training to enhance GPU memory efficiency. Despite these optimizations, memory usage and inference latency remain key concerns for real-world deployment. Future work could explore more efficient tokenization strategies to balance precision and computational efficiency. Second, our experiments were focused on retrospective EHR data, and its real-world applicability in prospective clinical settings remains to be validated. Further research could assess LabTOP’s performance in real-time hospital environments and evaluate its impact on clinical decision-making. Finally, generalization beyond ICU settings needs further investigation. Our current model is designed for short-term (near-future) prediction, leveraging high-density time-series data and trained in an auto-regressive manner. This setup aligns well with ICU and inpatient settings, where lab tests are recorded at high frequency and recent trends play a crucial role in clinical decision-making. In contrast, medical records for the general population tend to be recorded at irregular intervals, often spanning months or years between visits. Under such conditions, our training and inference strategy may not generalize effectively. Instead of directly applying LabTOP in this setting, a more suitable approach may involve alternative modeling techniques, such as point process-based models or learning strategies specifically designed for sparse and irregular medical data, which would be an interesting future research direction.
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (No.RS-2019-II190075), National Research Foundation of Korea (NRF) grant (NRF-2020H1D3A2A03100945, RS-2023-00262527), funded by the Korea government (MSIT).
@@ Appendix A Dataset Statistics
Time intervals for each lab item
Tables LABEL:tab:time_gaps_mimiciv, LABEL:tab:time_gaps_eicu and LABEL:tab:time_gaps_hirid present the time granularity (in hours) of lab items in MIMIC-IV, eICU, and HiRID, respectively. The time interval for each lab item represents the average interval (in hours) at which the lab item was recorded. A comparison of these tables reveals that the intervals between the same lab items are shortest in HiRID, indicating that lab tests are conducted more frequently in HiRID than in the other datasets.
tab:time_gaps_mimiciv Item Time Interval (Hours) Alanine Aminotransferase (ALT) 25.65 Albumin 40.22 Alkaline Phosphatase 25.79 Anion Gap 13.38 Aspartate Aminotransferase (AST) 25.62 Base Excess 7.63 Bicarbonate 13.34 Bilirubin, Total 25.53 Calcium, Total 14.09 Calculated Total CO2 7.63 Chloride 12.79 Creatinine 13.40 Free Calcium 10.33 Glucose 9.99 H 10.94 Hematocrit 13.26 Hemoglobin 12.80 I 10.94 INR (PT) 18.41 L 10.94 Lactate 10.98 Lactate Dehydrogenase (LD) 33.55 Lymphocytes 47.93 Magnesium 13.45 MCH 15.05 MCHC 15.05 MCV 15.05 Oxygen Saturation 9.34 PCO2 7.63 pH 7.85 Phosphate 14.00 Platelet Count 14.74 PO2 7.62 Potassium 12.43 Potassium, Whole Blood 9.93 PT 18.41 PTT 16.25 RDW 15.06 RDW-SD 14.41 Red Blood Cells 15.05 Sodium 12.42 Temperature 17.91 Urea Nitrogen 13.44 White Blood Cells 15.04 Average 15.80
tab:time_gaps_eicu Item Time Interval (Hours) Basos 25.80 Eos 25.78 Lymphs 24.57 Monos 24.60 Polys 24.85 Albumin 26.75 Alkaline Phos. 28.56 ALT (SGPT) 28.67 Anion Gap 17.54 AST (SGOT) 28.62 Base Excess 14.68 Bedside Glucose 4.10 Bicarbonate 17.44 BUN 17.57 Calcium 17.63 Chloride 17.29 Creatinine 17.53 FiO2 14.65 Glucose 16.64 HCO3 14.13 Hct 17.34 Hgb 16.82 Magnesium 20.06 MCH 20.21 MCHC 20.19 MCV 20.18 MPV 20.36 O2 Sat (%) 14.34 PaCO2 13.89 PaO2 13.91 pH 13.87 Phosphate 21.55 Platelets × 1000 19.97 Potassium 14.10 PT - INR 24.72 RBC 20.24 RDW 20.31 Sodium 15.43 Total Bilirubin 28.46 Total Protein 28.62 WBC × 1000 20.24 Average 19.81
tab:time_gaps_hirid Item Time Interval (Hours) Base Excess in Arterial Blood by Calculation 6.59 Bicarbonate [moles/volume] in Arterial Blood 6.59 C-Reactive Protein [mass/volume] in Serum or Plasma 24.72 Calcium, Ionized [moles/volume] in Blood 6.59 Carbon Dioxide [partial pressure] in Arterial Blood 6.41 Carboxyhemoglobin/Hemoglobin Total in Arterial Blood 6.58 Chloride [moles/volume] in Blood 6.64 Creatinine [moles/volume] in Blood 23.74 Glucose [moles/volume] in Serum or Plasma 3.85 Hemoglobin [mass/volume] in Arterial Blood 6.39 Hemoglobin [mass/volume] in Blood 17.67 INR in Blood by Coagulation Assay 21.09 Lactate [mass/volume] in Arterial Blood 6.68 Leukocytes [#/volume] in Blood 18.26 Magnesium [moles/volume] in Blood 26.61 MCH [entitic mass] 18.35 MCHC [mass/volume] in Cord Blood 18.35 MCV [entitic volume] 18.32 Methemoglobin/Hemoglobin Total in Arterial Blood 6.57 Oxygen [partial pressure] in Arterial Blood 6.39 Oxygen Saturation in Arterial Blood 6.57 pH of Arterial Blood 6.31 Phosphate [moles/volume] in Blood 24.72 Platelets [#/volume] in Blood 18.74 Potassium [moles/volume] in Blood 6.82 Sodium [moles/volume] in Blood 6.77 Urea [moles/volume] in Venous Blood 24.43 Average 12.99
@@ Appendix B Implementation Details
To create samples for these LLMs555Specifically, at the time of the experiments, GPT-4o referred to gpt-4o-2024-11-20 and GPT-4o-mini corresponded to gpt-4o-mini-2024-07-18., we followed the same strategy with GenHPF. In other words, we grouped test samples that have the same prior medical history (i.e., the lab items that occurred at the same time), and the LLM was instructed to answer values of the target lab items based on their prior sequence of medical events as prompt. The prompt we used for lab test outcome prediction can be found in Figure 3.
tab:mae_mimiciv Item Count MAE SMAPE Unit Glucose 72887 mg/dL Sodium 60140 mEq/L Potassium 60131 mEq/L Ph 59069 units Chloride 59000 mEq/L Hemoglobin 57786 g/dL Bicarbonate 56749 mEq/L Hematocrit 56497 % Anion Gap 56445 mEq/L Urea Nitrogen 56406 mg/dL Creatinine 56086 mg/dL Magnesium 54857 mg/dL Phosphate 52445 mg/dL Calcium, Total 51933 mg/dL Platelet Count 51197 K/uL Po2 50431 mm Hg Base Excess 50431 mEq/L Pco2 50379 mm Hg Calculated Total Co2 50329 mEq/L White Blood Cells 50194 K/uL Red Blood Cells 50153 m/uL Mch 50122 pg Rdw 50109 % Mchc 50066 %, g/dL Mcv 50045 g/dL Ptt 37743 sec Inr ( Pt ) 34670 None Pt 34659 sec L 32700 U Lactate 31327 mmol/L H 30831 U Rdw - Sd 29442 fL Free Calcium 28758 mmol/L I 25471 U Oxygen Saturation 18239 % Potassium, Whole Blood 16293 mEq/L Asparate Aminotransferase ( Ast ) 14980 IU/L Alkaline Phosphatase 14696 IU/L Alanine Aminotransferase ( Alt ) 14582 IU/L Bilirubin, Total 13902 mg/dL Temperature 10331 °C Lactate Dehydrogenase ( Ld ) 8792 IU/L Albumin 7552 g/dL Lymphocytes 7488 %
tab:mae_eicu Item Count MAE SMAPE Unit Bedside Glucose 179312 mg/dL Potassium 77648 mmol/L Sodium 71184 mmol/L Glucose 67292 mg/dL Hgb 65241 g/dL Chloride 64786 mmol/L Creatinine 64094 mg/dL Hct 63676 % Bun 63539 mg/dL Calcium 61862 mg/dL Bicarbonate 61027 mmol/L Platelets X 1000 56096 K/mcL Wbc X 1000 55922 K/mcL Rbc 55748 M/mcL Mcv 54438 fL Mchc 54131 g/dL Mch 51772 pg Rdw 51522 % Anion Gap 50748 nan Magnesium 38416 mg/dL Mpv 37782 fL Pao2 29405 mm Hg Paco2 29181 mm Hg Ph 28965 nan Monos 28891 % Lymphs 28774 % Hco3 27738 mmol/L Phosphate 26150 mg/dL Polys 25945 % Fio2 23978 % Albumin 23545 g/dL Base Excess 23264 mEq/L O2 Sat ( % ) 21639 % Ast ( Sgot ) 20840 Units/L Total Protein 20677 g/dL Alt ( Sgpt ) 20529 Units/L Alkaline Phos. 20391 Units/L Total Bilirubin 19399 mg/dL Pt - Inr 18298 ratio Eos 17160 % Basos 10767 %
tab:mae_hirid Item Count MAE SMAPE Unit Glucose [ Moles / Volume ] In Serum Or Plasma 45573 mmol/L Sodium [ Moles / Volume ] In Blood 24856 mmol/L Potassium [ Moles / Volume ] In Blood 24587 mmol/L Bicarbonate [ Moles / Volume ] In Arterial Blood 22615 mmol/L Base Excess In Arterial Blood By Calculation 22582 mmol/L Calcium. Ionized [ Moles / Volume ] In Blood 22396 mmol/L Carboxyhemoglobin / Hemoglobin. Total In Arterial Blood 22389 % Methemoglobin / Hemoglobin. Total In Arterial Blood 22346 % Lactate [ Mass / Volume ] In Arterial Blood 22321 mmol/L Chloride [ Moles / Volume ] In Blood 21994 mmol/L Oxygen Saturation In Arterial Blood 21781 % Hemoglobin [ Mass / Volume ] In Arterial Blood 20697 g/L Ph Of Arterial Blood 20673 None Oxygen [ Partial Pressure ] In Arterial Blood 20586 mmHg Carbon Dioxide [ Partial Pressure ] In Arterial Blood 20313 mmHg Hemoglobin [ Mass / Volume ] In Blood 10373 g/L Leukocytes [ # / Volume ] In Blood 10124 G/L Mchc [ Mass / Volume ] In Cord Blood 10037 g/L Mcv [ Entitic Volume ] 10016 fL Mch [ Entitic Mass ] 9992 pg Platelets [ # / Volume ] In Blood 9794 G/L Creatinine [ Moles / Volume ] In Blood 7424 umol/L C Reactive Protein [ Mass / Volume ] In Serum or Plasma 6834 mg/L Urea [ Moles / Volume ] In Venous Blood 6248 mmol/L Inr In Blood By Coagulation Assay 6010 None Phosphate [ Moles / Volume ] In Blood 5385 mmol/L Magnesium [ Moles / Volume ] In Blood 5195 mmol/L
Appendix C Additional Experimental Results
Performances by the time interval to the last measured time for the target lab test
To further illustrate the performance of our models and Naive model from another perspective, we conduct an additional analysis by grouping test set samples based on the time interval between the last measured time and the prediction time for the target lab item. Specifically, we categorized samples by whether the time interval to the last measured time for the target lab item is within 24 hours or not, and measured performances within each group. The results are shown in Table LABEL:tab:time_gaps_naive_labtop_mimiciv.
We observe that when the time interval between the last lab measurement and the prediction time is more than 24 hours, the performance gap between LabTOP and Naive becomes more significant. Since the Naive model directly copies the last observed value as the prediction, its performance remains relatively strong when lab values are measured frequently and do not change dynamically over short time spans. However, in longer time interval cases, where patient conditions may have changed significantly, the Naive model struggles while LabTOP maintains higher predictive accuracy (SMAPE). This result suggests that more computationally advanced methods like LabTOP are particularly valuable in clinically meaningful scenarios, where lab tests are infrequent and accurate forecasting is essential for early detection and intervention. Therefore, while the computational cost of LabTOP is higher, this trade-off is necessary to achieve reliable performance in challenging clinical cases.
tab:time_gaps_naive_labtop_mimiciv 24h 24h NMAE Naive 0.072 0.076 LabTOP SMAPE (%) Naive 15.91 20.39 LabTOP
LabTOP performances for individual lab items
The experimental results of LabTOP for each individual lab item in MIMIC-IV, eICU, and HiRID are presented in Table LABEL:tab:mae_mimiciv, Table LABEL:tab:mae_eicu, and Table LABEL:tab:mae_hirid, respectively.
tab:app_normal_abnormal Normal Abnormal NMAE Naive Naive() GenHPF XGBoost GPT-4o GPT-4o-mini LLaMA-3.1-Inst. LabTOP SMAPE (%) Naive Naive() GenHPF XGBoost GPT-4o GPT-4o-mini LLaMA-3.1-Inst. LabTOP
Model performances for normal and abnormal range values in MIMIC-IV
We group each lab test sample in the test split into normal and abnormal ranges using reference ranges provided in MIMIC-IV, and evaluate models separately for each group. These results are provided in Table LABEL:tab:app_normal_abnormal.
Exploration of different model configurations
We evaluate LabTOP’s performance at different model scales by varying the number of Transformer decoder layers, attention heads, and hidden dimension size. Specifically, we experiment with configurations comparable to BERT-mini, BERT-small, and BERT-base to analyze the effect of model size. The results of these experiments are presented in Table LABEL:tab:model_size.
tab:model_size Model Layers Attention Heads Hidden Dimension Parameter Size NMAE SMAPE LabTOP-mini 4 4 256 11M LabTOP-small 4 8 512 30M LabTOP-base 12 8 512 53M LabTOP-large 12 12 768 110M
Attention Map Analysis
To gain deeper insights into the prediction process of LabTOP, we can analyze an attention map to see how the model weighs different past clinical events when predicting a specific lab value. Here, since attention scores are generated for each predicted token across multiple layers and heads, these scores are systematically aggregated to construct a comprehensive attention representation for lab test outcome prediction. For each layer, the mean attention scores across all heads are computed, and these head-level averaged scores are then used to calculate the mean attention score for each layer. The aggregated attention score at each layer level is defined as:
where represents the final aggregated attention score, is the total number of layers, is the total number of attention heads, and denotes the attention score of head in layer . By averaging the attention maps computed for each predicted token, the final aggregated attention score is obtained. This resulting attention map captures attention scores from all preceding tokens before the target lab event.
To analyze attention at the event level, tokens before the target lab event are grouped based on their corresponding events. The total attention score for each event is computed by summing the attention scores of all tokens associated with that event Event-level attention scores are then visualized to determine which past events are most influential in predicting the lab test outcome.
Figure 4 illustrates the aggregated attention map during the prediction of a glucose outcome. The glucose level in the preceding events is recorded as . The model assigned a relatively high attention score to insulin-regular, which is the key factor in predicting a subsequent decrease in glucose levels. Since both the model prediction and actual glucose measurement confirmed this decrease, the analysis validates that the model effectively leverages relevant prior events to generate accurate predictions.